40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Aug 31 13:22:29 2012, Elli, Update, IOO, Correcting improper termination of the 55MHz input to the EOM 
    Reply  Fri Aug 31 18:32:47 2012, Koji, Update, IOO, Correcting improper termination of the 55MHz input to the EOM 
Message ID: 7332     Entry time: Fri Aug 31 18:32:47 2012     In reply to: 7328
Author: Koji 
Type: Update 
Category: IOO 
Subject: Correcting improper termination of the 55MHz input to the EOM 

 

The power switch of the frequency source was mechanically broken. I replace it with the spare in Kiwamu's treasure box.
I think this replaced new one also had the same problem before, we basically have to find an alternative switch.
The problem is too much friction between the metal housing and the plastic button. I have lubricated the button with
bicycle oil.

I have checked that the 55MHz output still have the output of 25dBm after a 5m BNC cable. (nominal: 27dBm)

I could not test it with the MC as the chambers have already been closed with the light doors.

ELOG V3.1.3-